Flawless Cosmetic Company manufactures and distributes several different products. The company currently uses a plantwide allocation method for allocating overhead at a rate of $6 per direct labor hour. Loren is the department manager of the Makeup Department which produces Products- Concealer (C) and Glow Cream (GC). Jennifer is the department manager of the Hair Care Department which manufactures Product Shampoo (S). The product costs (per case of 24 bottles) and other information are as follows: Products GC 115.00 $ 70.00 $42.00 11.00 12.00 195.00 $118.50 $65.00 3 600 Direct materials Direct labor Overhead 48.00 32.00 31.50 17.00 Machine hours 4 300 2 Number of cases (per year) 500 Based on this information, if Flawless changes its allocation basis to machine hours, what is the overhead rate per machine hour? o $3.00 O $2.00 O $17.00 O $6.325
Answer #1

Solution 1:

Total Overhead = number of cases* Overhead per case of all the 3 products

= (32*300)+ (17*500)+ (12*600)

= 9600 + 8500 + 7200

= $25,300

Total Machine Hours = (4*300) + (2*500) +(3*600) = 1200 + 1000 + 1800 = 4000 Mchine hours

Overhead rate per machine hour = Total Overhead / Total Machine hours = $25300 / 4000 = $6.325

Hence Last option is correct.

Solution 2:

Total Overhead = number of cases* Overhead per case of all the 3 products

= (24*300)+ (21*500)+ (13*600)

= 7200 + 10500 + 7800

= $25,500

Total Machine Hours = (4*300) + (2*500) +(3*600) = 1200 + 1000 + 1800 = 4000 Mchine hours

Overhead rate per machine hour = Total Overhead / Total Machine hours = $25500 / 4000 = $6.375 per machine hour

Total product cost per case for Product GC = $72+ $33 + ($6.375*2) = $117.75

Hence last option is correct.